OpenStudy (anonymous):

Medal You had a credit card balance of $378 before you used the card to purchase books for $160, CDs for $22, and a pair of shoes for $58. You then make a payment to the credit card company of $378. Find your updated balance after the payment is processed. $ ?

OpenStudy (haleyelizabeth2017):

What is 160+22+58?

OpenStudy (anonymous):

240

OpenStudy (haleyelizabeth2017):

Okay so what is 378-240?

OpenStudy (anonymous):

138

OpenStudy (haleyelizabeth2017):

Okay now 138+378?

OpenStudy (anonymous):

516

OpenStudy (danjs):

Updated Balance = (Previous Balance) + (Purchases) - (Payments)

OpenStudy (haleyelizabeth2017):

Good! :)

OpenStudy (anonymous):

so the answer is $516

OpenStudy (haleyelizabeth2017):

Wait @DanJS....wouldn't it be -purchases because that is spending money?

OpenStudy (mathstudent55):

The balance was $378 and a payment of $378 was made, so the new balance is zero before the new charges. Now just add the three new charges. That is the new balance.

OpenStudy (danjs):

No , purchases add to your balance that you owe, payments subtract from your debt

OpenStudy (haleyelizabeth2017):

oh!

OpenStudy (mathstudent55):

The new balance is: $160 + $22 + $58

OpenStudy (danjs):

Updated Balance = (Previous Balance) + (Purchases) - (Payments) = 378 + (160+22+58) - 378 = what @mathstudent55 said

OpenStudy (mathstudent55):

\(= \color{red}{\cancel{378}} + (160+22+58) - \color{red}{\cancel{378}}\) The two amounts in red cancel each other leaving only the new charges.

OpenStudy (anonymous):

240